OpenStudy (anonymous):

You invest $1,300 in an account that has an annual interest rate of 5%, compunded anually. How much money will be in the account after 10 years?

OpenStudy (anonymous):

Rounded to the nearest dollar would this be $1,950?

jimthompson5910 (jim_thompson5910):

Now you're using the formula A = P(1+r/n)^(nt) In this case, P = 1300, r = 0.05, n = 1 and t = 10 So you'll get A = 1300(1+0.05/1)^(1*10)

jimthompson5910 (jim_thompson5910):

No it should be a bit higher than 1950

OpenStudy (anonymous):

Hmmm, I cant seem to figure out what I did wrong? =( Can you tell me what I should be getting?

jimthompson5910 (jim_thompson5910):

let me write it out, one sec

jimthompson5910 (jim_thompson5910):

A = 1300(1+0.05/1)^(1*10) A = 1300(1+0.05/1)^(10) A = 1300(1+0.05)^(10) A = 1300(1.05)^(10) A = 1300*1.62889462677744 A = 2117.56301481068 A = 2118

OpenStudy (anonymous):

Thank you sooo much!! =)

jimthompson5910 (jim_thompson5910):

you're very welcome